# A heterotic integrable deformation of the principal chiral model

### v1

@@ 1 Introduction
One of the main impacts of the study of integrable model in high energy physics is the connection of quantities of integrable Green-Schwarz superstring $\sigma$-models in supercoset Anti de Sitter backgrounds and those of dual integrable conformal field theories [2, 3] in context of the AdS/CFT-duality. In the last years, some effort has been spent to generalise this to other, less symmetric settings. On the world-sheet side of the duality, new integrable models were found that are deformations of such (semi)symmetric (super)cosets, namely Yang-Baxter deformations [4, 5, 6], $\lambda$-[7] and WZ-term [8] deformations and their interplay [9], or generalisations to non-symmetric cosets [10, 11, 12]. So far, this survey was focussing on conventional two-dimensional non-linear $\sigma$-models, i.e. the embedding of a two-dimensional world-sheet in some $d$-dimensional target space. [^1] See [11, 13] for two recent reviews and more references.
@@
Similar to that, O$(d,d)$ generalised geometry appeared first in the Hamiltonian formulation and the analysis of the current algebra of the world-sheet string [33, 34, 35, 36] before it was considered as a tool for supergravity [37, 38, 39, 40, 41, 42, 43]. For the heterotic string the progress was reverse. Heterotic supergravity was phrased in terms of $O(d,d+n)$ generalised geometry [44, 45, 46, 43], but the role of O$(d,d+n)$ has not been illuminated for the heterotic world-sheet model until recently [1]. This is the latest addition to the program of duality covariant formulations of world-volume theories either via actions [47, 48, 49, 50, 51, 52] or via Hamiltonian formulations [53, 54, 55, 56, 57, 58, 59, 60]. Following [1], a heterotic $\mathcal{E}$-model is defined in section 2.
@@ 2 The classical bosonic heterotic string in the Hamiltonian formalism
For the purpose of this paper the generalised flux formulation is better suited. To this end, we diagonalise the generalised metric $\mathcal{H}_{\mathcal{A}\mathcal{B}}={E_{\mathcal{A}}}^{\mathcal{M}}{E_{\mathcal{B}}}^{\mathcal{N}}\mathcal{H}_{\mathcal{M}\mathcal{N}}$ with a generalised vielbein ${E_{\mathcal{A}}}^{\mathcal{M}}$ and some suitable constant $\mathcal{H}_{\mathcal{A}\mathcal{B}}$. The coupling to the background is encoded in the Poisson brackets via the so-called generalised fluxes $F_{\mathcal{A}\mathcal{B}\mathcal{C}}$:
@@
with $F_{\mathcal{A}\mathcal{B}\mathcal{C}}=3\partial_{\mathcal{[A}}{E_{\mathcal{B}}}^{\mathcal{N}}{E_{\mathcal{C]N}}}$. In the particular case that these $F_{\mathcal{ABC}}$ can be chosen to be constant, we would call this construction heterotic $\mathcal{E}$-model, in analogy to the construction for ordinary $2d$ $\sigma$-models and O$(d,d)$-generalised geometry. For a generic heterotic supergravity background these fluxes are not constant.
@@
00 0 $=\partial_{+}j_{-a}+\partial_{-}j_{+a}-2{F^{\beta}}_{ca}k_{\beta}j_{-}^{c}-H_{abc}j_{+}^{b}j_{-}^{c}$
00 0 $=\partial_{+}j_{-}^{a}-\partial_{-}j_{+}^{a}+{f^{a}}_{bc}j_{+}^{b}j_{-}^{c}$ (2.10)
00 0 $=2\partial_{-}k^{\alpha}-{F^{\alpha}}_{bc}j_{+}^{b}j_{-}^{c}$
@@ 3.1 Integrability of the undeformed model
This is close to the twist function form $\mathcal{R}(\lambda,\mu)=\mathcal{R}^{(0)}(\lambda,\mu)\phi^{-1}(\lambda)$ with a solution of the classical Yang-Baxter equation $\mathcal{R}^{(0)}$, like $\mathcal{R}(\lambda,\mu)=\frac{t_{a}\otimes t^{a}}{\lambda-\mu}$ and a twist function $\phi(\mu)$. This form ensures that this is a solution of the classical Yang-Baxter equation (3.6). Although not quite of this form, it is nevertheless easy to see that (3.7) solves the classical Yang-Baxter equation, as $\mathfrak{t}_{\alpha}\otimes t^{\alpha}$ is a central element of $\tilde{\mathfrak{g}}\otimes\tilde{\mathfrak{g}}$, and all its contributions drop out of the classical Yang-Baxter equation (3.6).
@@ 4 Discussion
  • Different starting point for the deformation:
@@
      non-abelian chiral currents $k^{\alpha}$, i.e. some non-trivial gauge algebra $h$ instead of $\mathfrak{u}(1)^{n}$
@@
Beside this, there are some questions that are not addressed in this article and left for future research:
Non-deformation limit.
As discussed briefly in the main text, the objects of the Lax construction, i.e. the Lax pair and the classical $\mathcal{R}$-matrix do not have the correct non-deformation limit. The $\mathfrak{g}$-valued part of the Lax connection (3.18) correctly reproduces the one of the principal chiral, but the full $\mathfrak{g}\oplus\mathfrak{u}(n)^{n})$-valued Lax connection does not reproduce the full undeformed one (3.3). A similar problem also arises for the poles and zeros of the twist function. Whereas $\phi_{\epsilon,h}\rightarrow\phi_{PCM}$ and both have double poles at 00 and infinity, $\phi_{\epsilon,h}=(1+\lambda)f_{\epsilon,h}(\lambda)\lambda^{-2}$ has three zeros:
@@
$\lambda_{1}=-1,\qquad\lambda_{2,3}=-\frac{1}{2\epsilon^{2}}\left(2\epsilon^{2}+(1+h)\pm\sqrt{8\epsilon^{2}+(h+1)^{2}}\right)$
@@
$\lambda_{2}=-\frac{1+h_{0}}{\epsilon^{2}}-\frac{3}{1+h_{0}}+\mathcal{O}(\epsilon^{2})\overset{\epsilon,h_{0}\rightarrow 0}{\longrightarrow}-\infty,\qquad\lambda_{3}=\frac{1}{1+h_{0}}+\mathcal{O}(\epsilon^{2})\overset{\epsilon,h_{0}\rightarrow 0}{\longrightarrow}=1$
@@
It would be interesting to see whether a different construction could be found which has a correct non-deformation limit.
Reduction of a conventional $\sigma$-model.
From the point of view of the systematics of integrable models, the question is: how new or exotic is this model really? The present Lax representation is valued in $\mathfrak{g}\oplus\mathfrak{u}(1)^{n}$. A natural candidate for a more conventional integrable $\sigma$-model, that would reproduce the same equations of motion (2.10), would be something like a principal chiral model or a WZW-model with target space $G\times T^{n}$, together with some constraint that makes $k_{-}$ (which did not appear in the construction here) vanish.
This is clearly possible for the undeformed model. For the deformed model this does not seem so simple and deserves further attention.
Supergravity and $\alpha^{\prime}$.
The usual question for the integrable deformations and their application in string theory is, whether the deformed backgrounds are also supergravity solutions, given that the undeformed one is. A detailed discussion is left for future research when deformations are known that start at a more realistic starting point than the principal chiral model. [^4] In comparison to the type II $\sigma$-models, one could for example not make a principal chiral model to a string $\sigma$-model by simply choosing a supergroup with vanishing Killing form.
Another connected issue when comparing to heterotic supergravity is the following. Matching the present parametrisation (2.6) to the standard one (i.e. [44, 1]) that would lead to the heterotic supergravity action, one would need to make the identification $A\rightarrow\sqrt{\alpha^{\prime}}A$ in (2.6). One notices that deformation parameter enters at exactly the same level as $\sqrt{\alpha^{\prime}}$. Hence, in cases where the starting point is a supergravity background, one might identify
$\epsilon^{2}\sim\alpha^{\prime},$ (4.2)
up to dimensionality, and the integrable deformation becomes the $\alpha^{\prime}$-correction.
@@ Acknowledgements
The author thanks Falk Hassler and Eric Lescano for comments on the draft.

### v2

@@ 1 Introduction
One of the main impacts of the study of integrable model in high energy physics is the connection of quantities of integrable Green-Schwarz superstring $\sigma$-models in supercoset Anti de Sitter backgrounds and those of dual integrable conformal field theories [2, 3] in context of the AdS/CFT-duality. In the last years, some effort has been spent to generalise this to other, less symmetric settings. On the world-sheet side of the duality, new integrable models were found that are deformations of such (semi)symmetric (super)cosets, namely Yang-Baxter deformations [4, 5, 6], $\lambda$-[7] and WZ-term [8] deformations and their interplay [9], or generalisations to non-symmetric cosets [10, 11, 12]. So far, this survey was focussing on conventional two-dimensional non-linear $\sigma$-models, i.e. the embedding of a two-dimensional world-sheet in some $d$-dimensional target space. See [11, 13] for two recent reviews.
@@
Similar to that, O$(d,d)$ generalised geometry appeared first in the Hamiltonian formulation and the analysis of the current algebra of the world-sheet string [33, 34, 35, 36] before it was considered as a tool for supergravity [37, 38, 39, 40, 41, 42, 43].
For the heterotic string the duality group is O$(d,d+n)$ [44, 45]. Heterotic supergravity was phrased in terms of $O(d,d+n)$ generalised geometry [46, 47, 48, 43]. World-sheet formulations of the heterotic string related to O$(d,d+n)$ generalised geometry have been suggested in [49], and recently [1]. This is the latest addition to the program of duality covariant formulations of world-volume theories either via actions [50, 51, 52, 53, 54, 55] or via Hamiltonian formulations [56, 57, 58, 59, 60, 61, 62, 63]. Following [1], a heterotic $\mathcal{E}$-model is defined in section 2.
@@ 2 The classical bosonic heterotic string in the Hamiltonian formalism
For the purpose of this paper the generalised flux formulation is better suited. To this end, we diagonalise the generalised metric $\mathcal{H}_{\mathcal{A}\mathcal{B}}={E_{\mathcal{A}}}^{\mathcal{M}}{E_{\mathcal{B}}}^{\mathcal{N}}\mathcal{H}_{\mathcal{M}\mathcal{N}}$ with a generalised vielbein ${E_{\mathcal{A}}}^{\mathcal{M}}$ and some suitable constant $\mathcal{H}_{\mathcal{A}\mathcal{B}}$. The coupling to the background is encoded in the Poisson brackets of the currents $J_{\mathcal{A}}={E_{\mathcal{A}}}^{\mathcal{M}}\mathcal{P}_{\mathcal{M}}$:
@@
via the so-called generalised fluxes $F_{\mathcal{A}\mathcal{B}\mathcal{C}}=3\partial_{\mathcal{[A}}{E_{\mathcal{B}}}^{\mathcal{N}}{E_{\mathcal{C]N}}}$. In the particular case that these $F_{\mathcal{ABC}}$ can be chosen to be constant, we would call this construction heterotic $\mathcal{E}$-model, in analogy to the construction for ordinary $2d$ $\sigma$-models and O$(d,d)$-generalised geometry. For a generic heterotic supergravity background these fluxes are not constant.
@@
00 0 $=\partial_{+}j_{-a}+\partial_{-}j_{+a}-2{F^{\beta}}_{ca}k_{\beta}j_{-}^{c}-H_{abc}j_{+}^{b}j_{-}^{c},$
00 0 $=\partial_{+}j_{-}^{a}-\partial_{-}j_{+}^{a}+{f^{a}}_{bc}j_{+}^{b}j_{-}^{c},$ (2.10)
00 0 $=2\partial_{-}k^{\alpha}-{F^{\alpha}}_{bc}j_{+}^{b}j_{-}^{c},$
@@ 3.1 Integrability of the undeformed model
This is close to the twist function form $\mathcal{R}(\lambda,\mu)=\mathcal{R}^{(0)}(\lambda,\mu)\phi^{-1}(\lambda)$ with a solution of the classical Yang-Baxter equation $\mathcal{R}^{(0)}$, like $\mathcal{R}(\lambda,\mu)=\frac{t_{a}\otimes t^{a}}{\lambda-\mu}$ and a twist function $\phi(\mu)$. This form is a sufficient condition in order for $\mathcal{R}$ to be a solution of the classical Yang-Baxter equation (3.6) and a typical assumption in the classification of integrable $\sigma$-models [66]. From this point of view it is noteworthy that this simple model has no $\mathcal{R}$-matrix of this form. Nevertheless, it is easy to see that (3.7) solves the classical Yang-Baxter equation, as $\mathfrak{t}_{\alpha}\otimes t^{\alpha}$ is a central element of $\tilde{\mathfrak{g}}\otimes\tilde{\mathfrak{g}}$, and all its contributions drop out of the classical Yang-Baxter equation (3.6).
@@ 4 Discussion
  • Different starting points for the deformation:
@@
      non-abelian chiral currents $k^{\alpha}$, i.e. some non-trivial gauge algebra $\mathfrak{h}$ instead of $\mathfrak{u}(1)^{n}$
@@
  • the deformed models as reduction of known integrable $\sigma$-models.
    From the point of view of the systematics of integrable models, the question is: how new is this model really? The present Lax representation is valued in $\mathfrak{g}\oplus\mathfrak{u}(1)^{n}$. A natural candidate for a more conventional integrable $\sigma$-model, that would reproduce the same equations of motion (2.10), would be something like a principal chiral model or a WZW-model with target space $G\times T^{n}$, together with some constraint that makes $k_{-}$ (which did not appear in the construction here) vanish. This is clearly possible for the undeformed model, but does not seem possible straightforward for the deformed model. This deserves further attention.
    A related question would be whether these models can be obtained from 4d Chern-Simons theory [20, 21, 22, 23, 24, 25]. The not quite ’twist function’-form of the $\mathcal{R}$-matrices (3.7) and (3.19) might suggest that this is not the case.
Beside these questions, two issues should be addressed and will be briefly discussed below – the questions about the embedding of these models into string theory and their incorrect non-deformation limit. Both questions certainly require more study, in particular if the present approach can be extended to the above mentioned more non-trivial examples.
4.1 Embedding into heterotic supergravity
The usual question for the integrable models and their deformations in context of string theory is, whether they are also supergravity backgrounds. A detailed discussion is left for future research when a similar analysis to the one in this article has been done with the inclusion of fermions.
This question consists of two parts: 1) When is the undeformed model a solution of heterotic supergravity. 2) Is the $A$-deformation a solution to heterotic supergravity, given the undeformed model is?
The principal chiral model.
In comparison to the type II $\sigma$-models, the heterotic principal chiral model cannot be completed a string $\sigma$-model by simply choosing a supergroup with vanishing Killing form [67].
The WZW-type undeformed model.
A valid starting point for the deformation in heterotic supergravity would be the WZW-type model, i.e. target space geometry corresponding to the fluxes (3.11) with $h=1$, $\epsilon=0$. This is clearly a solution of heterotic supergravity, as it only contains NS-NS fields and is a conformal model. WZW-models have been applied in several circumstances in heterotic string theory, for example for the heterotic string on AdS3 [68] or compactifications on group manifolds with torsion [69]. Some of these discussions actually require (chiral) gauged WZW-models [70, 71, 72]. For this, the present approach has to be extended from group to coset manifolds as discussed above.
The deformed model.
One can check that an integrable deformation of such a WZW-model will correspond to a solution of heterotic supergravity, given that the two deformation parameters $h$ and $\epsilon$ in (3.11) are related in a certain way.
For example, one of the equations of motion of heterotic supergravity [16] – in the frame formalism and for constant dilaton – is:
$0=R^{(+)}_{ab}-2\alpha^{\prime}\left(\mathrm{tr}(F_{ac}{F_{b}}^{c})-R^{(+)}_{ac}{{R^{(+)}}_{b}}^{c}\right).$ (4.2)
Here, $R^{(+)}$ is, as usual, the Ricci tensor of the torsionful connection $\Gamma+\frac{1}{2}H$, where $\Gamma$ is the Christoffel connection. The choice of fluxes (3.11) means that
$\mathrm{tr}(F_{ac}{F_{b}}^{c})$ $=\epsilon^{2}\kappa_{ab},\qquad R_{ab}^{(+)}=-\chi^{2}\kappa_{ab}$
with the Killing form $\kappa$ on $\mathfrak{g}$ and $\chi^{2}=\frac{1}{4}(h-1)^{2}(2+(h-1)^{2})$ for simplicity of notation.
So far, we ignored dimensions. Only for the following analysis, these are briefly reintroduced. In particular, $[\chi^{2}]=[\epsilon]=\left[\alpha^{\prime}\right]^{-1}$. The heterotic supergravity equation (4.2) at $\mathcal{O}(\alpha^{\prime})$ is satisfied, given that deformation parameters are related:
@@ 4.1 Embedding into heterotic supergravity
$\chi^{2}=\frac{1}{4\alpha^{\prime}}-\sqrt{\frac{1}{16\alpha^{\prime 2}}+\epsilon^{2}}\approx 2\alpha^{\prime}\epsilon^{2}+(\alpha^{\prime})^{-1}\mathcal{O}\left((\alpha^{\prime}\epsilon)^{4}\right).$ (4.3)
The other heterotic supergravity equations hold in a similar way. Also, by virtue of the Jacobi identity of $\mathfrak{g}$, the anomaly cancellation condition
$\mathrm{d}H=0=\mathrm{tr}(F\wedge F)$ (4.4)
is automatically fulfilled for fluxes of the deformed model (3.11).
4.2 Non-deformation limit.
As discussed briefly in the main text, the objects of the Lax construction, i.e. the Lax pair and the classical $\mathcal{R}$-matrix do not have the correct non-deformation limit. The $\mathfrak{g}$-valued part of the Lax connection (3.18) correctly reproduces the one of the principal chiral model, but the full $\mathfrak{g}\oplus\mathfrak{u}(n)^{n})$-valued Lax connection does not reproduce the full undeformed one (3.3). A similar problem also arises for the poles and zeros of the twist function. Whereas $\phi_{\epsilon,h}\rightarrow\phi_{PCM}$ and both have double poles at 00 and infinity, $\phi_{\epsilon,h}=(1+\lambda)f_{\epsilon,h}(\lambda)\lambda^{-2}$ has three zeros:
$\lambda_{1}=-1,\qquad\lambda_{2,3}=-\frac{1}{2\epsilon^{2}}\left(2\epsilon^{2}+(1+h)\pm\sqrt{8\epsilon^{2}+(h+1)^{2}}\right)$ (4.5)
@@
$\lambda_{2}=-\frac{1+h_{0}}{\epsilon^{2}}-\frac{3}{1+h_{0}}+\mathcal{O}(\epsilon^{2})\overset{\epsilon,h_{0}\rightarrow 0}{\longrightarrow}-\infty,\qquad\lambda_{3}=\frac{1}{1+h_{0}}+\mathcal{O}(\epsilon^{2})\overset{\epsilon,h_{0}\rightarrow 0}{\longrightarrow}=1$ (4.6)
@@
This result might seem puzzling, as for the known integrable deformations the non-deformation limit gives the expected undeformed Lax pairs, see e.g. [11]. An explanation for this might be the following: It is well-known in perturbation theory that deformations (or perturbations) of partial differential equations do not necessarily correspond to deformations of the solution space. One might interpret this result in this context, as the Lax pair corresponds to a solution of the system. Nevertheless, this question deserves further attention.
@@ Acknowledgements
The author thanks Falk Hassler, Eric Lescano and the anonymous referee for PRD for helpful comments on the draft.